OpenStudy (anonymous):

Fill in the blank. Sin ___________ equals √3/2 .

OpenStudy (anonymous):

60(?).. a little rusty on my trigonometry

OpenStudy (anonymous):

yup. definitely 60!

OpenStudy (anonymous):

thanx

OpenStudy (zarkon):

\[\frac{\pi}{3}+k\cdot2\pi\] \[k\] an integer

OpenStudy (zarkon):

and \[\frac{2\pi}{3}+k\cdot2\pi\]